# Security and Quality in LLM-Generated Code: A Multi-Language, Multi-Model Analysis

### v1

@@ 5 Results and Discussion
5.1 Compilation-Time Errors
The first part of Table V presents insights into the performance of different LLMs in generating compilable code without errors in four programming languages (Java, Python, C++, and C). To calculate the compilation success rate, let $P$ be the programming language, $C$ be the number of compilable files in $P$, $T$ be the total number of files from in $P$, and $S_{P}$ be the compilation success rate (in percentage). We then define $C=\sum_{i=1}^{n}C_{i}$, where $C_{i}$ is $1$ if file $i$ is compileable successfully and 00 otherwise. Moreover, we define $T=n$, where $n$ is the total number of files. We also define the compilation success rate in $P$ as $S_{P}=({C_{P}}/{T_{P}})\times 100$.
@@ 5.1 Compilation-Time Errors
Table V: Breakdown (%) of LLM-generated code files without compilation-time errors (error free) and semantic issues.
@@
The variability in the compilation-time error-free rates observed in Table V for Java, C++, and C suggests that LLMs exhibit different strengths depending on the language. For Java, performance varies more significantly compared to Python, with success rates ranging between 88.00% and 95.00%. Some models, such as claude-3.5 and GPT-4o, outperform others, while llama-3, gemini-1.5, and codestral show slightly lower success rates at 88%, 88.5%, and 88.5%, respectively.
These results can be attributed to Java’s verbose, statically typed nature, which reduces compilation-time errors when an LLM correctly understands its syntax and principles. The slight variations likely stem from differences in how well each LLM handles Java’s syntax rules, exceptions, and object-oriented features (i.e., encapsulation). Models like claude-3.5 and GPT-4o may be better fine-tuned for Java-specific dependencies and required imports, as missing import statements—such as java.utils.*—are a common issue affecting compilation success.
Takeaway. Models like claude-3.5 and GPT-4o perform better in Java, due to better handling of Java’s syntax rules and package dependencies, while models like llama-3, gemini-1.5, and codestral show lower success rates. This highlights the importance of fine-tuning for Java-specific aspects to improve LLM performance.
@@
Takeaway. LLMs suffer with C++ with missing include statements, incorrect type handling, different standards (i.e., C++11, C++17), and misinterpretation of APIs/functions. C++ syntax complexity, library usage, and type system contribute to frequent compile-time errors with STL and 3rd party APIs/libraries.
The varying success rates in C code generation are due to their ability in handling C programming standards and dependencies. The high success rate of claude-3.5 (96%) suggests that it effectively complied with C standards and included the necessary headers, avoiding common problems (i.e., undeclared types and missing libraries). Meanwhile, llama-3 with the lowest success rates (88. 00%), encountered frequent errors related to unknown types (i.e., bool), missing headers (i.e., cgi/cgi.h), and incorrect function arguments. These issues point to a reliance on non-standard libraries or incorrect assumptions about the development environment. These inconsistencies highlight the need to ensure that the generated code adheres to standard C practices, incorporates all required dependencies, and conforms to suitable function signatures to enhance the probability of successful compilation.
@@
Types of compilation-time errors. A breakdown of the compilation-time errors is summarized in Table VI. Python’s breakdown is excluded from the table since it lists only error types occurring more than once across all LLMs. In the following, compilation errors are grouped into categories to understand the common failure modes across the models.
  1. Library Errors The most common error category across all models was related to missing imports or incorrect package references. Errors such as ”cannot find symbol” and failure to recognize or include required third-party libraries, specificly missing import for java.utils.*, particularly for tasks that involved non-standard libraries, occurred frequently. This indicates a need for improved handling of Java libraries and packages, possibly by training LLMs on more diverse codebases with proper import statements. In addition, LLMs may benefit from an enhanced contextual understanding of external dependencies in programming tasks.
  2. Exception Handling Missing or incorrect exception handling was another area where the models faltered. gemini-1.5, codestral, and llama-3 were particularly prone to this type of error, indicating that these models could improve by focusing on the need to handle exceptions based on the exception that can be triggered by code in scope for a particular code block or method.
  3. Missing Class/Member An often recurrent issue flagged by codestral is the absence of a class or class member, such as setter and/or getter methods. This error arises when the model presumes the presence of a specific method or class without including the pertinent code or even suggesting it within comments.
  4. Syntax Syntax errors, including incorrect symbols, missing semicolons, or invalid syntax, were found primarily in codestral. This may suggest a need for fine-tuning on Java syntax conventions or a deeper understanding of language-specific rules.
  5. Type Compatibility Errors related to incompatible types, such as attempts to assign a java.lang.Object to a java.lang.String, or java.io.IOException cannot be converted to java.lang.SecurityException. These errors indicate the models’ ability to infer the data types in context.
  6. Variable Identifier Errors arising from the use of undefined variables were found in multiple models. This indicates that the models occasionally fail to maintain variable state consistency across different sections of the code.
@@
  8. Binding/Qualifier Errors related to failing to bind a variable correctly or mismatches with qualifiers. These errors often occur in function parameters or references.
  9. Constant Undeclared Errors with undeclared constants like true and nullptr.
Table VI: Breakdown of LLM-generated code files compilation-time error types. LLMs: ❶ claude-3.5, ❷ gemini-1.5, ❸ codestral, ❹ GPT-4o, ❺ llama-3.
@@
5.2 Semantic
The results shown in Table V under the semantic error-free part present the percentage of valid semantic code generated. The observed variance in success rates within the same LLM, with most models differing by approximately 8.50% in their ability to generate correct code in different languages, except gemini-1.5 which shows a range of 13.50%, highlights considerable variations in the way these models process the same coding task depending on the target programming language. Despite the same reasoning capabilities required, this variation suggests discrepancies in how each model interprets and implements programming logic, potentially due to differences in training data, model architecture, or optimization strategies. The wider range of gemini-1.5 may indicate particular challenges in understanding or generating code for certain languages or complex programming constructs, reflecting either gaps in its training data or special consideration in its design.
@@ 5.2 Semantic
Takeaway. The variance in success rates among different LLMs, with a range of 13.50% for gemini-1.5, indicates notable differences in how these models handle coding tasks, likely due to variations in training data, model design, and optimization strategies.
@@
For Java, the models generally showed strong performance, with claude-3.5 achieving the highest success rate of 95.00%. This indicates a good proficiency in creating effective and precise Java code, probably because of Java’s object-oriented design and consistent libraries, which apparently correspond well with the models’ training data. gemini-1.5 also performed remarkably well, reaching a success rate of 94.22%, demonstrating its proficiency in managing Java syntax and standard programming constructs. codestral and GPT-4o showed marginally lower success rates at 85.87% and 92.22%, respectively. The variation in performance, especially for codestral, could indicate difficulties in addressing Java’s more complex class structures or standard library usage, which aligns with the low success rate for the code compilation success rate as well. Consequently, the high success rates from all models indicate that these models are proficient in producing Java code, while also emphasizing the necessity for ongoing improvements to tackle the more complex elements of the language and reasoning capabilities.
Takeaway. LLMs perform well in the generation of Java code, with claude-3.5 leading at 95.00%, probably because of Java’s object-oriented design and consistent libraries. gemini-1.5 also excels, while codestral and GPT-4o show some challenges with complex structures. This highlights strong proficiency but also points to a need for improvement in handling Java’s complexities.
@@
Among the four languages evaluated, Python achieved the highest success rates, with gemini-1.5 at the top with a success rate of 97.34%, followed by claude-3.5 and GPT-4o with a success rate 96.72% and 96.35% respectively. Python’s simplicity and dynamic typing contribute to these high success rates, as its syntax is generally less strict and more flexible compared to statically typed languages. The high performance across all models in Python suggests they are proficient on the language and its constructs. Furthermore, llama-3, with a lower success rate of 91.56%, indicates that even models with generally high success rates can face challenges in generating correct Python code, potentially due to variations in handling language features and reasoning capabilities. codestral and GPT-4o demonstrated significant performance in generating functional C++ code, achieving success rates of 94.08% and 91.43%, respectively, reflecting effective handling of complex features in C++, followed by gemini-1.5 in third place with a success rate of 88.16% and llama-3 with the lowest success rate of 85.51%.
Takeaway. codestral and GPT-4o outperform at generating C++ code with success rates of 94.08% and 91.43%, effectively handling complex C++ features. Generation of C codes presents more challenges, with lower success rates in all models due to strict memory management and header file requirements. codestral leads in C at 88.97%, while llama-3 is lowest at 83.08%. These results highlight the need for improvements in handling C’s rigorous syntax and memory management demands.
The success rates in C were considerably lower for all models, with codestral achieving 88.97% and llama-3 reaching 83.08%. This lower performance can be attributed to strict C requirements for accurate memory management, header file inclusion, and model failures in solving a few tasks in C. claude-3.5 and GPT-4o performed slightly below codestral with success rates of 88.72% and 87.69% respectively. These results emphasize the need for targeted improvements in AI code generation to address the demanding syntax, header file inclusion, and memory management requirements.
5.3 Static Features
@@
This section evaluates the static code features of LLMs generated code across Java, Python, C++, and C using several complexity metrics, such as lines of code (LoC), cyclic complexity (CyC), cyclic complexity density (CCD) and cognitive complexity (CoC). The analysis included code generated by various LLMs identified as claude-3.5, gemini-1.5, codestral, GPT-4o, and llama-3. The results are provided in Table VII.
For Java code, as summarized in Table VII, shows that claude-3.5 produced the highest LoC (6,480) and the highest CoC (891). gemini-1.5 showed the highest CCD (0.23), indicating higher complexity. Conversely, codestral generated relatively less complex code with the lowest CyC (730) and CoC (585).
The evaluation of the Python code showed that among the models, gemini-1.5 had the highest CCD (0.37), reflecting a very dense complexity despite having a lower LoC (2,077). claude-3.5 produced the highest LoC (3,474) and the highest CoC (881) the same as gemini-1.5. codestral presented the lowest values in all metrics, indicating simpler and less complex code.
@@
Table VII: Static features: lines of code (LoC), cyclomatic complexity (CyC), density (CCD), and cognitive complexity (CoC).
@@ 5.4 Security and Quality Attribute
Table VIII presents Java results based on key quality attributes. For the LoC, claude-3.5 produces the highest LoC, increasing both complexity and variability. In contrast, codestral is the most efficient model regarding LoC. In terms of Security, which measures protection against unauthorized access and vulnerabilities, gemini-1.5 produces the most vulnerabilities among the LLMs for Java code. This result likely stems from a lack of security-focused training examples compared to other models.
Table VIII: Quality attributes against security (S), reliability (R), maintainability (M), and security hotspots (SH).
@@
According to Table VIII, codestral shows the fewest reliability issues, with a score of 37, indicating its ability to generate Java code with minimal crashes or errors. This may be linked to the lower line of code produced by codestral compared to other models. In terms of maintainability, codestral also performs best, with 535 issues, reflecting that its code is the easiest to understand, modify, and extend. This can be attributed to the use of clear structures, proper comments, and function decomposition. For security hotspots, llama-3 performs the best, with only 27 issues.
@@
For the Python code evaluation in Table VIII, the models show notable variations in quality. claude-3.5, generating 3,474 lines, has 82 maintainability issues and 44 security hotspots, indicating areas needing security improvements. codestral, with 2,077 lines, performs better with 50 maintainability issues and 39 security hotspots, reflecting better overall control. gemini-1.5 and GPT-4o, producing 3,064 and 2,906 lines respectively, show similar outcomes with maintainability issues (83 and 75) and security hotspots (36 and 43). llama-3, generating 2,633 lines, has the fewest security hotspots (34) and moderate maintainability issues (63), indicating balanced but not flawless performance.
@@
Table VIII reveals that claude-3.5, with 6,725 LoC, has the highest maintainability issues (590) but few security hotspots (36), suggesting strong security but complex maintenance. codestral, generating 4,293 LoC, has the fewest maintainability issues (470) and security hotspots (20), indicating a balanced performance. gemini-1.5, producing 5,822 LoC and with a higher count of both security issues (17) and maintainability concerns (587). GPT-4o and llama-3, with 5,263 and 5,077 lines, respectively, demonstrate moderate and balanced performance in both areas.
@@
For C,  Table VIII shows that claude-3.5, with 7,509 LoC, has a high number of maintainability issues (449) and security hotspots (186), indicating difficulties in managing large and secure code. codestral, producing 4,532 lines, has fewer security hotspots (115) and maintainability issues (314), though it scores lower in reliability (39). gemini-1.5, with 6,555 lines, shows the most security issues (33) and moderate maintainability concerns (352), signaling significant security problems. GPT-4o, generating 5,926 lines, balances moderate security hotspots (133) and maintainability issues (380), while llama-3, with 5,006 lines, shows higher reliability (54) but faces substantial security hotspots (182).
@@
CWE Categories of Security Quality Attribute. The detection of security flaws in the code produced by five Large Language Models (LLMs) for Java, Python, C, and C++ highlights substantial differences in both the quality and security of the code among the models and programming languages. The Table IX presents a breakdown of identified CWE categories, reflecting the different LLMs strengths and weaknesses in secure code generation. This analysis sheds light on common security challenges that various LLMs might present when producing code, identifying opportunities to enhance LLM-driven code generation.
@@
Table IX: Breakdown of the security quality attributes based on the CWE categories. LLMs: ❶ claude-3.5, ❷ gemini-1.5, ❸ codestral, ❹ GPT-4o, ❺ llama-3.
@@
5.5 Clean Code Attribute
Table X shows the clean code analysis for Java, where most models handled consistency well, with zero issues reported for all except GPT-4o (8) and llama-3 (61). However, intentionality issues were prominent, particularly for claude-3.5 (203) and gemini-1.5 (244), suggesting that the clarity of code purpose could be improved. Adaptability scores were highest for claude-3.5 (532) and lowest for codestral (182), indicating varying levels of code flexibility for future changes.

### v2

@@ 5 Results and Discussion
5.1 Semantic Soundness
Compilation-Time Errors. The first part of Table VIII presents insights into the performance of different LLMs in generating compilable code without errors in four programming languages (Java, Python, C++, and C). To calculate the compilation success rate, let $P$ be the programming language, $C$ be the number of compilable files in $P$, $T$ be the total number of files from in $P$, and $S_{P}$ be the compilation success rate (in percentage). We then define $C=\sum_{i=1}^{n}C_{i}$, where $C_{i}$ is $1$ if file $i$ is compileable successfully and 0 otherwise. Moreover, we define $T=n$, where $n$ is the total number of files. We also define the compilation success rate in $P$ as $S_{P}=({C_{P}}/{T_{P}})\times 100$.
@@ 5.1 Semantic Soundness
Table VIII: Breakdown (%) of LLM-generated code files without compilation-time errors (error free) and semantic issues.
@@
The variability in compilation-time error-free rates reported in Table VIII for Java, C++, and C indicates that LLMs exhibit different strengths depending on the programming language. For Java, performance fluctuates more than in Python, with success rates ranging from 88.00% to 95.00%. Models such as claude-3.5 and GPT-4o outperform others, while llama-3, gemini-1.5, and codestral show slightly lower rates at 88%, 88.5%, and 88.5%, respectively.
These results can be attributed to Java’s verbose, statically typed nature, which reduces compilation-time errors when an LLM correctly interprets its syntax and principles. The slight performance differences likely reflect how effectively each LLM handles Java’s syntax rules, exception handling, and object-oriented features (i.e., encapsulation). Models such as claude-3.5 and GPT-4o appear to be better fine-tuned for Java-specific dependencies and required imports. Missing import statements, such as java.util.*, remain a common cause of compilation failures.
Takeaway. Models like claude-3.5 and GPT-4o perform better on Java due to improved handling of syntax rules and package dependencies, while llama-3, gemini-1.5, and codestral show lower success rates, underscoring the importance of Java-specific fine-tuning to improve performance.
@@
Takeaway. LLMs struggle with C++ due to missing includes, type errors, standard mismatches (i.e., C++11 vs. C++17), and API misinterpretations. The language’s syntax complexity and strict type system lead to frequent compile-time errors, especially with STL and third-party libraries.
The varying success rates in C code generation reflect differences in handling C standards and dependencies. The high success rate of claude-3.5 (96%) indicates adherence to C conventions, including proper header inclusion and type declarations. In contrast, llama-3, with the lowest success rate (88.00%), frequently produced errors involving undeclared types (i.e., bool), missing headers (e.g., cgi/cgi.h), and incorrect function arguments. These issues suggest reliance on non-standard libraries or flawed assumptions about the compilation environment. Such inconsistencies highlight the need to generate code that follows standard C practices, includes necessary dependencies, and respects function signature conventions to improve compilation.
@@
Types of Compilation-Time Errors. A breakdown of the compilation-time errors is summarized in Table IX. Python’s breakdown is excluded from the table since it lists only error types occurring more than once across all LLMs. In the following, compilation errors are grouped into categories to understand the common failure modes across the models.
  1. Library Errors. The most common error category across all models involved missing imports or incorrect package references. Frequent issues included “cannot find symbol” errors and failures to recognize or include required third-party libraries, particularly missing imports such as java.util.*, especially in tasks involving non-standard libraries. These patterns suggest a need for better handling of Java libraries and packages, potentially through training LLMs on more diverse codebases with correct import usage. In addition, improving the contextual understanding of external dependencies may enhance model performance in such tasks.
  2. Exception Handling. Missing or incorrect exception handling was another area where the models faltered. gemini-1.5, codestral, and llama-3 were particularly prone to this type of error, indicating that these models could improve by focusing on the need to handle exceptions based on the exception that can be triggered by code in scope for a particular code block or method.
  3. Missing Class/Member. An often recurrent issue flagged by codestral is the absence of a class or class member, such as setter and/or getter methods. This error arises when the model presumes the presence of a specific method or class without including the pertinent code or even suggesting it within comments.
  4. Syntax Syntax errors, including incorrect symbols, missing semicolons, or invalid syntax, were found primarily in codestral. This pattern suggests a need for improved fine-tuning on Java syntax conventions and a deeper understanding of language-specific rules.
  5. Type Compatibility. Errors related to incompatible types, such as attempts to assign a java.lang.Object to a java.lang.String, or java.io.IOException cannot be converted to java.lang.SecurityException. These errors indicate the models’ ability to infer the data types in context.
  6. Variable Identifier. Errors arising from the use of undefined variables were found in multiple models. This indicates that the models occasionally fail to maintain variable state consistency across different sections of the code.
@@
  8. Binding/Qualifier. Errors related to failing to bind a variable correctly or mismatches with qualifiers. These errors often occur in function parameters or references.
  9. Constant Undeclared. Errors with undeclared constants like true and nullptr.
Table IX: Compilation-time error types in generated code. LLMs: ❶ claude-3.5, ❷ gemini-1.5, ❸ codestral, ❹ GPT-4o, ❺ llama-3.
@@
Semantic Correctness Using Unit Test. The results in Table VIII summarize semantic correctness across languages. Most models exhibit consistent performance, though accuracy varies with language complexity. claude-3.5 achieved up to 92.1%, while GPT-4o ranged from 90.9% (Python) to 81.1% (C). gemini-1.5 showed a broader spread, from 90.5% to 77.0%, and codestral ranged from 88.6% to 73.7%. llama-3 yielded the lowest scores overall, peaking at 88.5% in Python. These differences underscore the sensitivity of LLM outputs to both prompt formulation and target language. Although tasks require similar reasoning, performance discrepancies suggest underlying differences in how models interpret and implement logic, likely influenced by training data, architecture, or optimization strategies. The wider variability seen in gemini-1.5, llama-3, and codestral may reflect challenges in handling specific language features or complex constructs, pointing to design or training gaps.
@@
Takeaway. The variance in success rates between different LLMs, with an average of 14.7% for gemini-1.5, llama-3, and codestral, indicates notable differences in how these models handle coding tasks, likely due to variations in training data, model design, and optimization strategies.
@@
For Java, the models generally showed strong performance, with claude-3.5 achieving the highest success rate of 91.8%. This indicates a good proficiency in creating effective and precise Java code, probably because of Java’s object-oriented design and consistent libraries, which apparently correspond well with the models’ training data. gemini-1.5 also performed well, reaching a success rate of 89.5%, demonstrating its proficiency in managing Java syntax and standard programming constructs. codestral and llama-3 showed slightly lower success rates at 82.9% and 81.6%, respectively. The variation in performance, especially for codestral and llama-3, could indicate difficulties in addressing Java’s more complex class structures or standard library usage, which aligns with its relatively lower compilation success rate. On the whole, the high semantic correctness across models confirms their capacity to generate functional Java code, while also highlighting areas where deeper reasoning or structural precision remains necessary.
Takeaway. LLMs perform strongly on Java, with claude-3.5 achieving the highest score at 91.8%, likely reflecting Java’s structured object-oriented design and stable libraries. gemini-1.5 follows at 89.5%, while codestral and llama-3 score lower at 82.9% and 81.6%, respectively. These results indicate overall proficiency but also expose remaining challenges in handling Java-specific complexity.
@@
Among the evaluated languages, Python achieved the highest success rates, with claude-3.5 leading at 92.1%, followed by GPT-4o and gemini-1.5 at 90.9% and 90.5%, respectively. Python’s simplicity and dynamic typing likely contribute to these high success rates, as its syntax is generally more permissive and flexible than statically typed languages. The consistently strong performance across models suggests they are well-versed in Python constructs. Notably, llama-3 achieved a slightly lower success rate of 88.5%, indicating that even high-performing models may encounter challenges in generating correct Python code, potentially due to differences in language handling or reasoning strategies. For C++, claude-3.5 achieved the highest semantic success rate at 91.2%, demonstrating a strong capability in handling C++ syntax and object-oriented constructs. GPT-4o followed at 88.0%, showing reliable performance on complex language features. codestral reached 83.7%, while gemini-1.5 and llama-3 trailed at 81.0% and 79.3%, respectively, indicating more pronounced challenges in generating correct C++ code across these models. Although these trends, such as Python demonstrating fewer compilation errors than C++, align with practitioners’ expectations, our results quantify the magnitude of these differences across models and link them to specific language quality and security outcomes.
Takeaway. claude-3.5 achieves the highest success rate in C++ at 91.2%, then GPT-4o at 88.0% and codestral at 83.7%, indicating differences in handling C++ complexity. C remains more challenging across models due to stricter syntax, with claude-3.5 leading at 84.2% and llama-3 scoring lowest at 72.9%. These results underscore ongoing limitations in low-level language reasoning and generation.
Success rates in C were notably lower across all models, with claude-3.5 achieving the highest at 84.2%, followed by GPT-4o at 81.1%, gemini-1.5 at 77.0%, codestral at 73.7%, and llama-3 at 72.9%. This performance gap can be attributed to C’s stricter requirements, including manual memory management, explicit type declarations, and the mandatory inclusion of standard headers. These findings underscore the need for targeted enhancements in LLM-based code generation to better address the syntactic rigor and low-level control demanded by the C programming language.
Takeaway. Semantic correctness is high in Python and Java for top-performing models, but drops in C++ and especially C, where success rates range roughly from 73% to 84%. This gap shows that the benchmark is not uniformly easy and exposes language-dependent challenges, particularly in lower-level languages with manual memory management and stricter type systems.
5.2 Complexity
@@
This section evaluates the static code features of LLMs generated code across Java, Python, C++, and C using several complexity metrics, such as lines of code (LoC), cyclic complexity (CyC), cyclic complexity density (CCD) and cognitive complexity (CoC). The analysis included code generated by various LLMs identified as claude-3.5, gemini-1.5, codestral, GPT-4o, and llama-3. The results are provided in Table X. For Java code, as summarized in Table X, shows that claude-3.5 produced the highest LoC (6,480) and the highest CoC (891). gemini-1.5 showed the highest CCD (0.23), indicating higher complexity. Conversely, codestral generated relatively less complex code with the lowest CyC (730) and CoC (585). The evaluation of the Python code showed that among the models, gemini-1.5 had the highest CCD (0.37), reflecting a very dense complexity despite having a lower LoC (2,077). claude-3.5 produced the highest LoC (3,474) and the highest CoC (881) the same as gemini-1.5. codestral presented the lowest values in all metrics, indicating simpler and less complex code.
@@
Table X: Static features: lines of code (LoC), cyclomatic complexity (CyC), density (CCD), and cognitive complexity (CoC).
@@ 5.3 Security and Quality Attribute
Table XI presents Java results across key quality attributes. For lines of code (LoC), claude-3.5 generates the highest output, contributing to increased complexity and variability. In contrast, codestral is the most efficient model in terms of LoC. Regarding security, which assesses resistance to unauthorized access and vulnerabilities, gemini-1.5 yields the highest number of vulnerabilities among the LLMs for Java. This outcome is likely due to a relative lack of security-focused training data compared to other models.
Table XI: Quality attributes against security (S), reliability (R), maintainability (M), and security hotspots (SH).
@@
Per Table XI, codestral has the fewest reliability issues (37), indicating minimal crashes or errors, likely due to its lower lines of code compared to other models. It also ranks best in maintainability, with 535 issues, reflecting code that is easier to understand, modify, and extend, attributable to clear structure, comments, and function decomposition. For security hotspots, llama-3 performs best, with only 27 issues.
@@
For the Python code evaluation in Table XI, the models show notable variations in quality. claude-3.5, generating 3,474 lines, has 82 maintainability issues and 44 security hotspots, indicating areas needing security improvements. codestral, with 2,077 lines, performs better with 50 maintainability issues and 39 security hotspots, reflecting better overall control. gemini-1.5 and GPT-4o, producing 3,064 and 2,906 lines respectively, show similar outcomes with maintainability issues (83 and 75) and security hotspots (36 and 43). llama-3, generating 2,633 lines, has the fewest security hotspots (34) and moderate maintainability issues (63), indicating balanced but not flawless performance.
@@
Table XI reveals that claude-3.5, with 6,725 LoC, has the highest maintainability issues (590) but few security hotspots (36), suggesting strong security but complex maintenance. codestral, generating 4,293 LoC, has the fewest maintainability issues (470) and security hotspots (20), indicating a balanced performance. gemini-1.5, producing 5,822 LoC and with a higher count of both security issues (17) and maintainability concerns (587). GPT-4o and llama-3, with 5,263 and 5,077 lines, respectively, demonstrate moderate and balanced performance in both areas.
@@
For C, Table XI shows that claude-3.5, with 7,509 LoC, has a high number of maintainability issues (449) and security hotspots (186), indicating difficulties in managing large and secure code. codestral, producing 4,532 lines, has fewer security hotspots (115) and maintainability issues (314), though it scores lower in reliability (39). gemini-1.5, with 6,555 lines, shows the most security issues (33) and moderate maintainability concerns (352), signaling significant security problems. GPT-4o, generating 5,926 lines, balances moderate security hotspots (133) and maintainability issues (380), while llama-3, with 5,006 lines, shows higher reliability (54) but faces substantial security hotspots (182).
To illustrate our findings, Appendix C presents examples of reliability, maintainability, and security hotspot issues in LLM-generated code. These cases show how seemingly minor flaws such as missing input validation, outdated API usage, or unsafe regular expressions can compromise code quality, even in relatively simple tasks.
@@
CWE Categories of Security Quality Attribute. The detection of security flaws in the code produced by five Large Language Models (LLMs) for Java, Python, C, and C++ highlights substantial differences in both the quality and security of the code among the models and programming languages. The Table XII presents a breakdown of identified CWE categories, reflecting the different LLMs strengths and weaknesses in secure code generation. This analysis sheds light on common security challenges that various LLMs might present when producing code, identifying opportunities to enhance LLM-driven code generation.
@@
To illustrate identified security vulnerabilities, Appendix C presents two representative cases: RSA encryption without OAEP padding (CWE-780) and hard-coded passwords (CWE-259), corresponding to tasks 106 and 83. These examples show that LLMs may default to insecure practices when safer alternatives exist, indicating limited contextual security awareness. These CWEs are well known and documented in prior analyses of LLM-generated and human-written code. The goal here is to quantify how often such known weaknesses, along with other vulnerabilities, appear across recent LLMs and programming languages under a unified benchmark. The patterns in Table XII therefore support comparative analysis of model and language behavior rather than claims of novel vulnerability types.
Table XII: Breakdown of the security quality attributes based on the CWE categories. LLMs: ❶ claude-3.5, ❷ gemini-1.5, ❸ codestral, ❹ GPT-4o, ❺ llama-3.
@@
5.4 Clean Code
Table XIII shows the clean code analysis for Java, where most models handled consistency well, with no issues reported except for GPT-4o (8) and llama-3 (61). However, intentionality issues were prominent, especially for claude-3.5 (203) and gemini-1.5 (244), suggesting that the clarity of code purpose could be improved. Adaptability scores were highest for claude-3.5 (532) and lowest for codestral (182), indicating varying levels of code flexibility for future changes.